Two point charges 40uc and 10uc are separated by a distance of 0.9m in air. Find the position of point where resistant intensity is zero.


1
Expert's answer
2021-03-14T19:17:25-0400

Let the distance between point charges "q_1"and "q_2" be "r". Let the position where the resultant electric field intensity is zero located at a distance "d" from the "q_1=40\ \mu C" charge and distance "(r-d)" from the "q_2=10\ \mu C" charge. Then, we can find this position by equating electric fields due to point charges "q_1"and "q_2" and solving for "d" :


"E_1=E_2,""\dfrac{kq_1}{d^2}=\dfrac{kq_2}{(r-d)^2},""\dfrac{d^2}{(r-d)^2}=\dfrac{q_1}{q_2},""\dfrac{d}{r-d}=\sqrt{\dfrac{q_1}{q_2}},""d=\dfrac{r}{3}\cdot\sqrt{\dfrac{q_1}{q_2}},""d=\dfrac{0.9\ m}{3}\cdot\sqrt{\dfrac{40\cdot10^{-6}\ C}{10\cdot10^{-6}\ C}}=0.6\ m."

Therefore, the resultant electric field intensity is zero at 0.6 m away from the "40\ \mu C" point charge and 0.3 m away from the "10\ \mu C" point charge.